A package of mass 50 kg is dropped from a helicopter and experiences a drag force of 120 N due to air resistance. Find the resultant force and use it to calculate the acceleration of the package during the fall. Take the acceleration due to gravity as g = 9.81 m s-2 and give your answer in m s-2 to 1 decimal place.

Answer:    Answer m s-2
Answer #1

Solr: Given, Mass (m) - 50kq & Force(P): 120N * Take acceleration due to gravity cg) = 9.81 mlsz => kle know, Hewtons and la
Answer #2

Fw = mass x g = 50 kg x 9.81 m s-2 =  490.5 N


Fw = 490.5n         Fdrag = 120n


Resultant force = 490.5n - 120n = 370.5n


acceleration = 
force=
370.5m s-2  = 7.41m s-2
mass50

to 1 decimal place = 7.4m s-2